HOME ESR EXPERIMENTS LABS
E121 E127 E132 E125 E143 laser_cooling_2021 E142 128
  nuclear two-photon decay, Page 1 of 12  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  234   Sat Jul 3 08:12:22 2021 Dima, YuriGeneralBeam stopWe stopped the beam at 08:10

The latest recorded tdms files:
  
  233   Sat Jul 3 07:35:31 2021 The remote Canadian eager beaver (Iris)AnalysisCloser potential isomer tracksHere's a list of the spectra to look at for
the potential isomer, taken btw 21:00 on
July 2nd and 6:30 on July 3rd. Two examples
 410MHz-2021.07.02.21.42.36.139.tiq.png410MHz-2021.07.02.22.40.24.894.tiq.png 
  232   Sat Jul 3 06:55:57 2021 Dima, DavidDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000761
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003720
  
  230   Sat Jul 3 06:48:38 2021 The remote Canadian eager beavers (Guy & Iris)AnalysisCloser potential isomer tracksAfter looking through >800 shots from Jul
2nd 21h to Jul 3rd 05h, we couldn't find
a single example of a "smoking gun" isomer
 70Se_close_pot_isomers.png 
  229   Sat Jul 3 04:58:34 2021 Dima, DavidDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000735
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003589
  
  228   Sat Jul 3 04:33:50 2021 The remote Canadian eager beavers (Guy & Iris)AnalysisSlight shift in 70Se frequencyLooks like the beam is slightly drifting...the
70Se peak frequency is slightly shifted to
the left, by ~0.5 kHz compared to earlier
 70Se-shift-21h-4h.png 
  227   Sat Jul 3 03:41:48 2021 The remote Canadian eager beavers (Guy & Iris)AnalysisSummed Spectra SummaryTo summarise the method of bulk summing the
spectra, I've created frequency vs time spectra
by using the e143_analyser.py script to stack
 72Ge_sum.png72Br_sum.png70Se_not-scraped.png70Se_strong-scraped.png70Se_not-scraped_high_contrast.png 
  226   Sat Jul 3 03:07:43 2021 The remote Canadian eager beavers (Guy & Iris)AnalysisSome single spectra (410 MHz) from 21-24hAttached are some single spectra with a rough
attempt to group them according to the different
ions that appear.
 July2-70Se-single-spectra-21-24h.pdf 
  225   Sat Jul 3 02:56:04 2021 Dima, DavidDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000707
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003452
  
  224   Sat Jul 3 01:28:08 2021 The remote Canadian eager beavers (Guy and Iris)AnalysisSummed spectra 410 MHz from 21:00-23:59 (70Se)Here's a picture with all injections with
an ion in the spectrum (empty ones were removed-
list below). The time slices are 0-0.5-1-1.5-2s,
 July2-21-0h-410MHz_AllSpectra_70Se.png 
  223   Sat Jul 3 00:59:01 2021 Dima, DavidDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000680
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003322
  
  222   Fri Jul 2 23:09:22 2021 Dima, David, LilianaDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000655
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003199
  
  221   Fri Jul 2 21:07:31 2021 LilianaDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000627
 p:\E143\iq\IQ_2021-06-30_23-27-34\0003062
  
  220   Fri Jul 2 20:33:44 2021 YuriAccelerator70Se setting with closed scrapersWe tune back to 70Se.
The scrapers stay closed at EEXDS2HA=0 mm,
EEXDS1HG=+5 mm
  
  219   Fri Jul 2 20:01:16 2021 LilianaRunsSome plots from the Br setting   6x 
  218   Fri Jul 2 19:57:12 2021 Shahab, Yuri, Wolfram, LilianaRunsstop DAQ At 19:56 the DAQ is stopped. 
The plan is to go back to the setting for
Se. 
  
  217   Fri Jul 2 19:06:18 2021 LilianaDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000600
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002931
  
  216   Fri Jul 2 17:31:04 2021 LilianaGeneralstatusat 17:20 intensity was suddenly reduced. 
Operators checked. Intensity was a bit low
and fluctuating for a while (8e8 to 1e9)
  
  215   Fri Jul 2 17:27:56 2021 LilianaDAQNTCAP statusThe latest recorded tdms files:
 p:\E143\sc\SC_2021-06-30_23-27-45\0000577
 p:\E143\iq\IQ_2021-06-30_23-27-34\0002818
  
ELOG V3.1.2-bd75964